### v2

@@ 1 Introduction
The Internet of Things (IoT) seamlessly integrates our cyber world with the physical world and is becoming increasingly prevalent in daily life. According to published statistics, the number of IoT devices has exceeded 15 billion and is projected to reach approximately 30 billion by 2030 [1].
@@
Despite the rapid proliferation of devices and providers, security remains a critical issue. IoT devices can be more challenging to secure than conventional devices due to their diverse hardware, software, and varied manufacturer profiles [2]. Moreover, unlike conventional devices, many IoT devices have nonstandard interfaces that impair user interaction and make it difficult to mitigate against vulnerabilities [3]. The NETSCOUT Threat Intelligence Report [4] indicates that a new IoT device on the network typically faces its first attack within 5 hours and becomes a specific attack target within 24 hours, and that most attacks exploit vulnerabilities in IoT devices [5]. Beyond being targets in attacks, compromised IoT devices can serve as tools in botnet attacks, exemplified by Mirai, URSNIF, and BASHLITE [6], where captured devices are used to orchestrate high-volume Distributed Denial of Service (DDoS) attacks. While various methods have been developed to detect and mitigate ongoing attacks, primarily Intrusion Detection Systems (IDS), proactive measures such as device-specific updates, internet access restrictions or isolation can prevent these attacks before they happen. Given the variety, quantity and unfamiliar interfaces of IoT devices, it is impractical to rely on users to implement these solutions. As a result, Device identification (DI) systems have been developed to automatically identify devices based on their activity and make it possible to apply appropriate security measures to create a secure IoT ecosystem. It is important to clarify that DI in this context aims to identify the device type (i.e., its make and model), and not to distinguish between unique instances of the same device type, which is a separate security challenge.
@@
Most existing approaches to DI involve the construction of Machine Learning (ML) models. When developing ML models, it is important to ensure that they generalize beyond the specific environment in which they were trained. In DI, this means that models trained on data from one device should be able to detect other devices of the same make and model operating within different network environments. Achieving this requires accounting for the domain shifts and variations typical of network and IoT environments [7]. In this regard, a prominent limitation of previous work is that most studies have relied upon a single dataset collected in a single network environment for both developing and testing their models [8, 9, 10]. This has resulted in an incomplete, and potentially misleading, picture of DI generalizability. Where multiple datasets have been used, the focus has been on the generalizability of methodologies and feature sets rather than of model instances [11, 12, 13].
@@
    Commitment to transparency and reproducibility: To foster transparency and encourage further research, we openly share our code and analysis [^1] Source code available at github.com/kahramankostas/GeMID, providing the community with the resources to replicate and build upon our findings.
@@ 2 Related Work
The field of IoT DI has been developing in earnest since 2017. Among the pioneering works is IoTSentinel [14], which builds models using 23 features extracted from packet headers. Packet headers, analogous to envelopes for data traveling over a network, contain information for routing and processing the data. By analyzing these headers, various useful features for network tasks can be extracted. Figure 1 illustrates a sample network packet and lists the header features present in it. Extractable features from packet headers include source/destination IP/MAC addresses, protocol, Time-to-Live (TTL), flag information, port numbers, sequence/acknowledgment numbers, and checksum. A significant contribution of this study was the introduction of  Aalto University IoT device captures [15], one of the most widely used open-access datasets in the IoT DI domain. Subsequent studies, such as IoTSense [9], expanded the feature sets by incorporating payload features (such as payload size and entropy) alongside header-derived features. Concurrently, other research focused on statistics derived from headers, emphasizing the relationship between packets within a specified range (e.g., TTL [8], packet size [16], time [16]).
@@
In IoTSentinel and IoTSense, features are derived from individual packet headers, but ML models are trained using features collected from multiple packets, known as fingerprints. These fingerprints consist of 12 packets in IoTSentinel and 20 packets in IoTSense. Despite using multiple packets, these fingerprints simply concatenate individual packet features, rather than generating inter-packet features. Alternatively, SysID [10] and IoTDevID [13] create fingerprints based on individual packet features, avoiding the need for merged data features. Obtaining fingerprints from an individual packet addresses the transfer problem that occurs when multiple devices share the same MAC address, a common issue due to the use of MAC addresses in the preassembly of packets [13].
@@
Deriving features from flow data instead of packet headers is also common. A flow refers to a sequence of packets sent from a specific source to a specific destination, considered as part of a single session or connection. The size of the flow can be a predetermined number of packets or packets within a specific time/protocol interval. These features can include basic flow information (source, destination, protocol), time-related metrics (start-end time, duration), packet and byte counts (total, average, max, min sizes), rate metrics (packet, byte, flow rate), flag statistics, inter-arrival times (average, max, min), and flow direction indicators (packets/bytes sent between source and destination). Sivanathan et al. [17] used features obtained from network flows, such as flow volume, flow duration, and average flow rate. They also introduced the widely used UNSW-DI dataset. Many other DI studies have also used flow features [18, 19].
@@
While flow-based features, such as flow volume and inter-arrival times, are prevalent [18, 19], they inherently encode network-specific traits—e.g., latency or bandwidth— rather than device characteristics. Studies like [17] and [20] exemplify this approach, yet their generalizability remains untested across diverse environments. Our analysis shows these methods are fragile, contrasting with packet-based features used in  [9, 14], which we validate as more reliable.
@@
Features can also be obtained from packet headers or flows through sliding windows of various sizes. For example, features such as port ranges, packet size, packet quantity, availability time and inter-arrival time can be subjected to statistical analysis within the window to derive features such as maximum, minimum, mean and standard deviation [20]. This inherently involves features based on multiple packets. However, an issue with features derived from flow/network statistics and windowing methods is that they are likely to capture implicit characteristics of the network environment, rather than just the characteristics of individual devices. In this paper, we argue that this makes them fragile, and demonstrate that they are not a reliable basis for carrying out DI.
@@
Another approach moves away from hand-crafted features by using deep learning to automate the feature engineering process. A common technique is to use the raw bytes of a network packet directly, as illustrated in Figure 1. In this method, the first n bytes are fed into a model—often a Convolutional Neural Network (CNN)—where each byte is treated as an image pixel [21, 22, 23, 24, 25]. This data is truncated or padded to a fixed length n. This concept extends beyond raw packets to other data representations. For instance, DeviceMien [12] uses stacked LSTM-Autoencoders on TCP-flow data, while Lopez-Martin et al. [26] employ a hybrid CNN-RNN model to treat traffic flows as pseudo-images. Others focus on temporal patterns, such as using a CNN to learn long-term relationships from short-term packet arrival sequences [27].
@@
However, using raw data for DI presents other problems that have not yet been adequately addressed. In studies where the entire raw packet data was used [22, 24], this included headers, payloads, and metadata, making it difficult to filter out identifying information such as MAC/IP addresses or string identifiers, both of which can significantly impair model generalizability. Whilst this dependency can be reduced by solely using packet payloads [23, 28], caution is needed because much of today’s internet traffic is encrypted. In practice encrypted payloads are likely to be ineffective for device identification models as the data appears random and opaque, providing no useful information about the contents. Thus, the added complexity of raw data is unlikely to contribute to the model’s performance and may even hinder it.
@@
One prominent approach focuses on the analysis of periodic background traffic, which is often characteristic of IoT device behavior. AUDI [29] and BehavIoT [30] exemplify this method. AUDI uses unsupervised learning, including Fourier transforms and signal autocorrelation, to model the periodic communication patterns of devices for identification. BehavIoT extends this by modeling not just periodic events but also user-initiated and aperiodic events to create comprehensive device and system behavior models from network traffic. These techniques capitalize on the insight that many IoT devices maintain constant, regular communication with backend servers, providing a stable basis for fingerprinting.
@@
Another technique involves identifying fine-grained, packet-level signatures. The PINGPONG [31] system demonstrates that short, unique sequences of packet lengths and directions can serve as reliable fingerprints for specific device events (e.g., a light turning on). These signatures often appear as request-reply pairs (“ping-pong”) between a device and a cloud server and can be extracted automatically from traffic without deep statistical analysis.
@@
Other methods have leveraged metadata from different network protocols. For instance, IoTFinder [32] uses passive DNS traffic, modeling a device’s identity based on the set of domain names it queries and the frequency of those queries. This DNS-based fingerprinting approach is designed to be scalable and effective even when devices are hidden behind Network Address Translation (NAT).
@@
While these diverse techniques have proven effective within their respective evaluation contexts, a critical challenge that many early studies did not fully address is the generalizability of the resulting models. Work by Kolcun et al. [33] highlighted this issue by demonstrating that the accuracy of various ML models for device identification degrades significantly over time when tested on data collected weeks or months after the training period, suggesting that static models are insufficient. Complementing this, Ahmed et al. [34] conducted a comprehensive analysis across multiple datasets, showing that temporal, spatial, and data-collection-methodology differences all impact fingerprinting accuracy, and that features robust in one context may not be in another. These findings underscore the fundamental problem of model generalization, which is the primary focus of our work.
@@
A related issue with existing DI studies is the lack of cross-dataset validation, which involves evaluating a model trained on one dataset against a different dataset to assess its generalizability and robustness. While some studies use multiple datasets [11, 12, 13], they typically apply their methods separately to each, training and testing within the same dataset (e.g., training on a portion of Dataset A and testing on the remainder). This approach raises potential concerns about their performance in unseen environments. It also exacerbates the specific issues raised above surrounding the use of statistical features, since models using these features would—in effect—be tested within the same network environment as the one they were trained in. In this study, we demonstrate the importance of cross-dataset validation in developing generalizable models.
@@
Many of the studies cited in this section, such as [9, 10, 11, 12, 13, 14, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 28, 27, 31, 33, 30, 29, 32, 26, 12, 34] validate their approaches using training and testing sets drawn from the same dataset or environment. Our work diverges by focusing on the stricter challenge of model instance generalizability, where a specific model trained on data from one environment must perform accurately on data from a completely different environment. This evaluation distinction makes a direct F1-score comparison with these works an ‘apples-to-oranges’ comparison and is the reason our study implements its own baselines to ensure a fair and controlled experiment.
@@ 3.1.1 UNSW
For feature and model selection, we utilized the IoT Attack Traces—ACM SOSR 2019 (UNSW-AD) [35] and IoT Traffic Traces—IEEE TMC 2018 (UNSW-DI) [17] datasets. The UNSW-DI dataset is tailored for DI tasks, comprising data from 24 benign devices collected over a 60-day period in 2016. In contrast, the UNSW-AD dataset, originally designed for anomaly detection (AD), contains 27 days of benign data and 17 days of both benign and malicious data from 28 devices collected over a 44-day period in 2018. For the UNSW-AD dataset, we used only the benign data.
@@
Figure 3: Devices represented in the MonIoTr dataset from both UK (blue) and USA (yellow) sites, along with their intersection (grey). Adapted from [36].
@@ 3.1.2 MonIoTr
After selecting the features and models, to remove any bias, we evaluate their effectiveness on a different dataset, IoT Information Exposure—IMC’19 (MonIoTr) [36]. The MonIoTr dataset consists of 81 devices. An important characteristic of this dataset is that data collection took place at two separate sites, one in the UK (MonIoTr-UK) and the other in the USA (MonIoTr-USA). Of the 81 devices, 26 devices were present in both sites, with the USA site having 47 devices (including 21 unique devices) and the UK site having 34 devices (including 8 unique devices). Figure 3 shows the distribution of devices by country.
@@
Many devices interact heavily with the local network, applications, tools, and cloud services during startup, so we merged power and interaction data into one active class. In a previous study  [37], we observed that when both active and idle data are available, active data is much more effective for DI. Consequently, we excluded idle data from this study.
@@ 4.1 Predictive Features
A voting system was used for feature selection, with each non-zero kappa value (with a tolerance of ±$\pm 5\%$, so $\geq 0.05$), indicating the feature had a positive contribution, resulting in a vote. Figure 6 illustrates this system. Features receiving four or more votes from either DD (green) or SS (red) categories, with at least one DD vote, advanced to the next stage. Features not meeting this criterion were eliminated. The requirement for at least one DD vote is based on its greater strictness in assessing generalizability. As a result of voting, we identified 46 of the 332 features as being individually predictive.
Figure 4: (Left) Data usage from UNSW datasets across different evaluation contexts. CV, SS, and DD steps are visualized in shades of blue, red, and green, respectively, throughout the paper for clarity. In the nomenclature, CV|AD-S1 denotes cross-validation on the AD dataset session 1. For cases other than CV, the first dataset is used for training and the second for testing. For example, AD-S1|DI-S2 means the first session of the UNSW-AD dataset is used for training and the second session of the UNSW-DI dataset is used for testing. (Right) The legend displayed is shared across Figures 5-7 for consistency and clarity.
Figure 5: Comparison of feature utility in UNSW datasets measured using CV (blue) and isolated methods, SS (red) and DD (green). CV tends to overestimate feature utility, with higher scores for many attributes. SS and DD produce more realistic evaluations. The discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models. For the legend, please refer to Figure 4.
Figure 6: Features identified as being predictive within one or more evaluation contexts (CV: blue, SS: red, DD: green), based on non-zero kappa scores of DT models. Note that the CV votes are not used to select features, but are shown here for completeness. For the legend, please refer to Figure 4.

### v3

@@ 1 Introduction
The Internet of Things (IoT) seamlessly integrates our cyber world with the physical world and is becoming increasingly prevalent in daily life. According to published statistics, the number of IoT devices has exceeded 15 billion and is projected to reach approximately 30 billion by 2030 [statista].
@@
Despite the rapid proliferation of devices and providers, security remains a critical issue. IoT devices can be more challenging to secure than conventional devices due to their diverse hardware, software, and varied manufacturer profiles [zarpelao2017survey]. Moreover, unlike conventional devices, many IoT devices have nonstandard interfaces that impair user interaction and make it difficult to mitigate against vulnerabilities [williams2017privacy]. The NETSCOUT Threat Intelligence Report [modi2019] indicates that a new IoT device on the network typically faces its first attack within 5 hours and becomes a specific attack target within 24 hours, and that most attacks exploit vulnerabilities in IoT devices [marzano2018evolution]. Beyond being targets in attacks, compromised IoT devices can serve as tools in botnet attacks, exemplified by Mirai, URSNIF, and BASHLITE [antonakakis2017understanding], where captured devices are used to orchestrate high-volume Distributed Denial of Service (DDoS) attacks. While various methods have been developed to detect and mitigate ongoing attacks, primarily Intrusion Detection Systems (IDS), proactive measures such as device-specific updates, internet access restrictions or isolation can prevent these attacks before they happen. Given the variety, quantity and unfamiliar interfaces of IoT devices, it is impractical to rely on users to implement these solutions. As a result, Device identification (DI) systems have been developed to automatically identify devices based on their activity and make it possible to apply appropriate security measures to create a secure IoT ecosystem. It is important to clarify that DI in this context aims to identify the device type (i.e., its make and model), and not to distinguish between unique instances of the same device type, which is a separate security challenge.
@@
Most existing approaches to DI involve the construction of Machine Learning (ML) models. When developing ML models, it is important to ensure that they generalize beyond the specific environment in which they were trained. In DI, this means that models trained on data from one device should be able to detect other devices of the same make and model operating within different network environments. Achieving this requires accounting for the domain shifts and variations typical of network and IoT environments [zhou2022domain]. In this regard, a prominent limitation of previous work is that most studies have relied upon a single dataset collected in a single network environment for both developing and testing their models [meidan2017detection, bezawada2018behavioral, aksoy2019automated]. This has resulted in an incomplete, and potentially misleading, picture of DI generalizability. Where multiple datasets have been used, the focus has been on the generalizability of methodologies and feature sets rather than of model instances [chowdhury2020network, ortiz2019devicemien, kostas22IoTDevID].
@@
    Commitment to transparency and reproducibility: To foster transparency and encourage further research, we openly share our code and analysis [^1] To maintain review anonymity, the source code link will be provided after the review process., providing the community with the resources to replicate and build upon our findings.
@@ 2 Related Work
The field of IoT DI has been developing in earnest since 2017. Among the pioneering works is IoTSentinel [miettinen2017iot], which builds models using 23 features extracted from packet headers. Packet headers, analogous to envelopes for data traveling over a network, contain information for routing and processing the data. By analyzing these headers, various useful features for network tasks can be extracted. Figure 1 illustrates a sample network packet and lists the header features present in it. Extractable features from packet headers include source/destination IP/MAC addresses, protocol, Time-to-Live (TTL), flag information, port numbers, sequence/acknowledgment numbers, and checksum. A significant contribution of this study was the introduction of  Aalto University IoT device captures [aalto2017dataset], one of the most widely used open-access datasets in the IoT DI domain. Subsequent studies, such as IoTSense [bezawada2018behavioral], expanded the feature sets by incorporating payload features (such as payload size and entropy) alongside header-derived features. Concurrently, other research focused on statistics derived from headers, emphasizing the relationship between packets within a specified range (e.g., TTL [meidan2017detection], packet size [kawai2017identification], time [kawai2017identification]).
@@
In IoTSentinel and IoTSense, features are derived from individual packet headers, but ML models are trained using features collected from multiple packets, known as fingerprints. These fingerprints consist of 12 packets in IoTSentinel and 20 packets in IoTSense. Despite using multiple packets, these fingerprints simply concatenate individual packet features, rather than generating inter-packet features. Alternatively, SysID [aksoy2019automated] and IoTDevID [kostas22IoTDevID] create fingerprints based on individual packet features, avoiding the need for merged data features. Obtaining fingerprints from an individual packet addresses the transfer problem that occurs when multiple devices share the same MAC address, a common issue due to the use of MAC addresses in the preassembly of packets [kostas22IoTDevID].
@@
Deriving features from flow data instead of packet headers is also common. A flow refers to a sequence of packets sent from a specific source to a specific destination, considered as part of a single session or connection. The size of the flow can be a predetermined number of packets or packets within a specific time/protocol interval. These features can include basic flow information (source, destination, protocol), time-related metrics (start-end time, duration), packet and byte counts (total, average, max, min sizes), rate metrics (packet, byte, flow rate), flag statistics, inter-arrival times (average, max, min), and flow direction indicators (packets/bytes sent between source and destination). Sivanathan et al. [sivanathan2018classifying] used features obtained from network flows, such as flow volume, flow duration, and average flow rate. They also introduced the widely used UNSW-DI dataset. Many other DI studies have also used flow features [kostas23thesis, jmila2022survey].
@@
While flow-based features, such as flow volume and inter-arrival times, are prevalent [kostas23thesis, jmila2022survey], they inherently encode network-specific traits—e.g., latency or bandwidth— rather than device characteristics. Studies like [sivanathan2018classifying] and [msadek2019iot] exemplify this approach, yet their generalizability remains untested across diverse environments. Our analysis shows these methods are fragile, contrasting with packet-based features used in  [bezawada2018behavioral, miettinen2017iot], which we validate as more reliable.
@@
Features can also be obtained from packet headers or flows through sliding windows of various sizes. For example, features such as port ranges, packet size, packet quantity, availability time and inter-arrival time can be subjected to statistical analysis within the window to derive features such as maximum, minimum, mean and standard deviation [msadek2019iot]. This inherently involves features based on multiple packets. However, an issue with features derived from flow/network statistics and windowing methods is that they are likely to capture implicit characteristics of the network environment, rather than just the characteristics of individual devices. In this paper, we argue that this makes them fragile, and demonstrate that they are not a reliable basis for carrying out DI.
@@
Another approach moves away from hand-crafted features by using deep learning to automate the feature engineering process. A common technique is to use the raw bytes of a network packet directly, as illustrated in Figure 1. In this method, the first n bytes are fed into a model—often a Convolutional Neural Network (CNN)—where each byte is treated as an image pixel [kotak2021iot, hoang2022data, hu2021robust, liu2022autonomous, yao2022cnn]. This data is truncated or padded to a fixed length n. This concept extends beyond raw packets to other data representations. For instance, DeviceMien [ortiz2019devicemien] uses stacked LSTM-Autoencoders on TCP-flow data, while Lopez-Martin et al. [lopez2017network] employ a hybrid CNN-RNN model to treat traffic flows as pseudo-images. Others focus on temporal patterns, such as using a CNN to learn long-term relationships from short-term packet arrival sequences [ma2020pinpointing].
@@
However, using raw data for DI presents other problems that have not yet been adequately addressed. In studies where the entire raw packet data was used [hoang2022data, liu2022autonomous], this included headers, payloads, and metadata, making it difficult to filter out identifying information such as MAC/IP addresses or string identifiers, both of which can significantly impair model generalizability. Whilst this dependency can be reduced by solely using packet payloads [hu2021robust, kotak2022iot], caution is needed because much of today’s internet traffic is encrypted. In practice encrypted payloads are likely to be ineffective for device identification models as the data appears random and opaque, providing no useful information about the contents. Thus, the added complexity of raw data is unlikely to contribute to the model’s performance and may even hinder it.
@@
One prominent approach focuses on the analysis of periodic background traffic, which is often characteristic of IoT device behavior. AUDI [marchal2019audi] and BehavIoT [hu2023behaviot] exemplify this method. AUDI uses unsupervised learning, including Fourier transforms and signal autocorrelation, to model the periodic communication patterns of devices for identification. BehavIoT extends this by modeling not just periodic events but also user-initiated and aperiodic events to create comprehensive device and system behavior models from network traffic. These techniques capitalize on the insight that many IoT devices maintain constant, regular communication with backend servers, providing a stable basis for fingerprinting.
@@
Another technique involves identifying fine-grained, packet-level signatures. The PINGPONG [trimananda2019pingpong] system demonstrates that short, unique sequences of packet lengths and directions can serve as reliable fingerprints for specific device events (e.g., a light turning on). These signatures often appear as request-reply pairs (“ping-pong”) between a device and a cloud server and can be extracted automatically from traffic without deep statistical analysis.
@@
Other methods have leveraged metadata from different network protocols. For instance, IoTFinder [perdisci2020iotfinder] uses passive DNS traffic, modeling a device’s identity based on the set of domain names it queries and the frequency of those queries. This DNS-based fingerprinting approach is designed to be scalable and effective even when devices are hidden behind Network Address Translation (NAT).
@@
While these diverse techniques have proven effective within their respective evaluation contexts, a critical challenge that many early studies did not fully address is the generalizability of the resulting models. Work by Kolcun et al. [kolcun2021revisiting] highlighted this issue by demonstrating that the accuracy of various ML models for device identification degrades significantly over time when tested on data collected weeks or months after the training period, suggesting that static models are insufficient. Complementing this, Ahmed et al. [ahmed2022analyzing] conducted a comprehensive analysis across multiple datasets, showing that temporal, spatial, and data-collection-methodology differences all impact fingerprinting accuracy, and that features robust in one context may not be in another. These findings underscore the fundamental problem of model generalization, which is the primary focus of our work.
@@
A related issue with existing DI studies is the lack of cross-dataset validation, which involves evaluating a model trained on one dataset against a different dataset to assess its generalizability and robustness. While some studies use multiple datasets [chowdhury2020network, ortiz2019devicemien, kostas22IoTDevID], they typically apply their methods separately to each, training and testing within the same dataset (e.g., training on a portion of Dataset A and testing on the remainder). This approach raises potential concerns about their performance in unseen environments. It also exacerbates the specific issues raised above surrounding the use of statistical features, since models using these features would—in effect—be tested within the same network environment as the one they were trained in. In this study, we demonstrate the importance of cross-dataset validation in developing generalizable models.
@@
Many of the studies cited in this section, such as [bezawada2018behavioral, aksoy2019automated, chowdhury2020network, ortiz2019devicemien, kostas22IoTDevID, miettinen2017iot, kawai2017identification, sivanathan2018classifying, kostas23thesis, jmila2022survey, msadek2019iot, kotak2021iot, hoang2022data, hu2021robust, liu2022autonomous, yao2022cnn, kotak2022iot, ma2020pinpointing, trimananda2019pingpong, kolcun2021revisiting, hu2023behaviot, marchal2019audi, perdisci2020iotfinder, lopez2017network, ortiz2019devicemien, ahmed2022analyzing] validate their approaches using training and testing sets drawn from the same dataset or environment. Our work diverges by focusing on the stricter challenge of model instance generalizability, where a specific model trained on data from one environment must perform accurately on data from a completely different environment. This evaluation distinction makes a direct F1-score comparison with these works an ‘apples-to-oranges’ comparison and is the reason our study implements its own baselines to ensure a fair and controlled experiment.
@@ 3.1.1 UNSW
For feature and model selection, we utilized the IoT Attack Traces—ACM SOSR 2019 (UNSW-AD) [hamza2019detecting] and IoT Traffic Traces—IEEE TMC 2018 (UNSW-DI) [sivanathan2018classifying] datasets. The UNSW-DI dataset is tailored for DI tasks, comprising data from 24 benign devices collected over a 60-day period in 2016. In contrast, the UNSW-AD dataset, originally designed for anomaly detection (AD), contains 27 days of benign data and 17 days of both benign and malicious data from 28 devices collected over a 44-day period in 2018. For the UNSW-AD dataset, we used only the benign data.
@@
Figure 3: Devices represented in the MonIoTr dataset from both UK (blue) and USA (yellow) sites, along with their intersection (grey). Adapted from [ren-imc19].
@@ 3.1.2 MonIoTr
After selecting the features and models, to remove any bias, we evaluate their effectiveness on a different dataset, IoT Information Exposure—IMC’19 (MonIoTr) [ren-imc19]. The MonIoTr dataset consists of 81 devices. An important characteristic of this dataset is that data collection took place at two separate sites, one in the UK (MonIoTr-UK) and the other in the USA (MonIoTr-USA). Of the 81 devices, 26 devices were present in both sites, with the USA site having 47 devices (including 21 unique devices) and the UK site having 34 devices (including 8 unique devices). Figure 3 shows the distribution of devices by country.
@@
Many devices interact heavily with the local network, applications, tools, and cloud services during startup, so we merged power and interaction data into one active class. In a previous study  [kostas2023externally], we observed that when both active and idle data are available, active data is much more effective for DI. Consequently, we excluded idle data from this study.
@@ 4.1 Predictive Features
A voting system was used for feature selection, with each non-zero kappa value (with a tolerance of $\pm 5\%$, so $\geq 0.05$), indicating the feature had a positive contribution, resulting in a vote. Figure 6 illustrates this system. Features receiving four or more votes from either DD (green) or SS (red) categories, with at least one DD vote, advanced to the next stage. Features not meeting this criterion were eliminated. The requirement for at least one DD vote is based on its greater strictness in assessing generalizability. As a result of voting, we identified 46 of the 332 features as being individually predictive.
Figure 4: (Left) Data usage from UNSW datasets across different evaluation contexts. CV, $S$, and $D$ steps are visualized in shades of blue, red, and green, respectively, throughout the paper for clarity. In the nomenclature, CV|AD-S1 denotes cross-validation on the AD dataset session 1. For cases other than CV, the first dataset is used for training and the second for testing. For example, AD-S1|DI-S2 means the first session of the UNSW-AD dataset is used for training and the second session of the UNSW-DI dataset is used for testing. (Right) The legend displayed is shared across Figures 5-7 for consistency and clarity.
Figure 5: Comparison of feature utility in UNSW datasets measured using CV (blue) and isolated methods, SS (red) and $D$ (green). CV tends to overestimate feature utility, with higher scores for many attributes. SS and $D$ produce more realistic evaluations. The discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models. For the legend, please refer to Figure 4.
Figure 6: Features identified as being predictive within one or more evaluation contexts (CV: blue, $S$: red, $D$: green), based on non-zero kappa scores of DT models. Note that the CV votes are not used to select features, but are shown here for completeness. For the legend, please refer to Figure 4.